OpenStudy (anonymous):

If p-1 is a factor of (p^4)+(p^2)+p-k. Then what is the value of K? I will give a medal!!

OpenStudy (anonymous):

@perl

OpenStudy (perl):

lets see

OpenStudy (perl):

Suppose we have the function f(p) = (p^4)+(p^2)+p-k if p -1 is a factor of (p^4)+(p^2)+p-k then f(1) = 0 so plug in 1

OpenStudy (perl):

not quite , i wrote f(1) = 0

OpenStudy (anonymous):

so 0=(1^4)+(1^2)+1-k

OpenStudy (perl):

yes that looks better

OpenStudy (anonymous):

so would it be K=3?

OpenStudy (anonymous):

Thank you @perl

OpenStudy (perl):

the intuition here is if i say p-1 is a factor of (p^4)+(p^2)+p-k that means (p^4)+(p^2)+p-k = (p - 1 ) * ( cubic expression in p ) then it makes sense that if p =1 , you will get zero ( 1-1) * ( something) = 0